If 22.9 grams of an aqueous solution of aluminum sulfate, Al2(SO4)3, contains 5.24 grams of aluminum sulfate, what is the percentage by mass of aluminum sulfate in the solution?________ % Al2(SO4)3
Answer #1

The formula to calculate mass percentage is :

Mass Percentage = ( Mass of solute / Mass of solution ) × 100%

( Solute is the substance which is present in lesser amount in a solution )

Here solute is aluminium sulfate Al2(SO4)3 which is dissolved in water to make an aqueous solution.

So mass of solute ( Al2(SO4)3) = 5.24 g

Mass of solution = 22.9 g

So using the formula of mass percentage, we get

Percentage by mass of Al2(SO4)3 = (Mass of Al2(SO4)3 / Mass of solution ) × 100 % = (5.24 g/ 22.9 g) × 100 % = (0.229) × 100% = 22.9% .